# Large Language Model for Table Processing: A Survey

### v2

@@ 5.1 Common Workflow of LLM-powered Agents
Research such as Chain-of-Thought (CoT) [20] and ReAct [69] prompt LLMs iteratively, organizing the reasoning process into multiple intermediate steps. Thus, LLMs address simpler subproblems step by step and progressively build a coherent response. Agent systems are designed to follow this paradigm and typically include modules such as memory, planning, and action [21]. The memory module stores state or observations from the environment or records past actions. This information can be utilized for future planning. The planning module chooses which action the agent needs to do in the current step, while the action module interacts with the environment and executes the action to get the outcomes. As shown in Figure 6, for table tasks, the agent first observes the table data and user intent. It generates prompts, decomposes complex tasks, plans actions, executes them on the table environment, and then updates the state or observation. This iterative process is repeated until expectations are met. Studies like SheetAgent [36], Chain-of-Table [29], ReAcTable [15] and $E^{5}$ [70] follow this workflow.
@@ 5.2.2 Complex Task Decomposition
Inspired by the CoT and least-to-most [71] prompting methods, researchers instruct LLMs to decompose complex table tasks into simpler sub-tasks. DIN-SQL [72] proposes breaking down the NL2SQL task into subtasks. The process involves identifying the relevant tables and columns associated with the query and producing intermediate sub-queries. For web tables, Dater [31] exploits LLMs as decomposers by breaking down huge evidence (a huge table) into sub-evidence (a small table) and decomposing a complex question into simpler sub-questions and getting SQLs.
@@ 5.3.2 Spreadsheet Manipulation and Data Analysis
Spreadsheet manipulation and data analysis are highly flexible. First, users’ expectations and requirements vary. Second, there is a wide range of operations and software APIs. Third, a complex task may contain multiple operations, leading to dynamic alterations in the table content.
@@ 5.3.3 Multiple Table Tasks
StructGPT [77] is capable of solving table QA, NL2SQL, and knowledge graph QA by developing three types of actions that target web tables, databases, and knowledge graphs. These actions function as a reader, extracting knowledge from the data, which is then used to assist the LLM in its future planning. The actions for web tables include extracting data or column names, whereas for databases, the actions involve extracting table data or metadata. UniDM [34] manages a variety of data manipulation tasks, such as data cleaning, on data lakes through a three-action process. The first action is extracting context information, including table metadata and related records associated with the task, utilizing this information as demonstrations or background knowledge. The second action involves converting this context information into NL texts suitable for LLMs to reason. The final action employs prompt engineering to formulate the desired prompt. TAP4LLM [78] is a pre-processing toolbox to generate table prompts. Its action includes: 1) selecting appropriate rows and columns from the table (called Table Sampling), 2) integrating them with other table data (relevant external knowledge, metadata) (called Table Augmentation), and 3) serializing this information to fit the context length of the LLM.
@@ 5.4.2 Role-play
5.5 Summary and Disscussion
The LLM-powered agent method combines the power of LLMs and the flexibility of external tools.
Cost Agents must prompt LLMs repeatedly to fulfill users’ expectations. Consequently, this increases the time and financial costs.
@@ 4 Excel online forums,
synthetically generated 1000k tables (png, HTML)
Traditional benchmarks, such as WikiTableQuestion [30], WikiSQL [82], and Spider [40], are already widely used and studied. We will not elaborate on those here but rather focus on new benchmarks. Table 2 presents recently proposed datasets and benchmarks, along with their data sources, sizes. We summarize the features of these new benchmarks as follows:
@@
  • Human Involved Labeling Some new datasets, which target data analysis, require extensive manual annotation. DS-1000 [43], InfiAgent-DABench [17], Tapilot-Crossing [80] are designed to assess data analysis tasks. They gather data from the internet and annotate it either automatically or semi-automatically. DS-1000 collects questions from StackOverflow, assesses their usefulness manually, and curates to form the benchmark. The authors manually adapt the original questions by providing input and output context into test cases and rewriting problems to prevent LLMs from learning and memorizing the data. InfiAgent-DABench invites human experts to evaluate the dataset quality and compare human-made and GPT-4 generated data analysis questions via multiple metrics. Tapilot-Crossing aims to build a benchmark for real-world data analysis.
  • Real-world Workload Most datasets are derived from the web or synthesized using templates. These publicly available online data are usually simple, as some of them are just tutorials for beginners. In contrast, tables in real-world scenarios are far more complex than these. SpreadsheetBench [37] proposes a benchmark aimed at real-world scenarios, where the authors meticulously analyzed user questions from four Excel forums. They argue that their benchmark could assess the performance of handling complex user instructions.
  • Larger Scale AnaMeta [87] is a large-scale table metadata dataset, and GitTables [89] downloads millions of CSV tables from GitHub and aligns them with knowledge bases. These two datasets can evaluate whether the AI system could understand the table schema and benchmark tasks like column type annotation. ComplexTable [44] consists of over 1,000k image tables, most of which have complex structures.
@@ 7 Challenges and Future Directions
  • Implicit User Intentions. When processing tables, users’ intentions are often unclear. For instance, in scenarios where the underlying tables are primarily stored in databases and hidden from end users, these users may have no knowledge of the underlying tables and columns. Consequently, they might issue vague or overly long queries with redundant information. Conversely, when the table content is provided by end users seeking analysis, they often pose general questions like “help me analyze the table” without a clear objective. In the former scenario, LLMs are encouraged to extract the specific intention from the verbose query. In the latter scenario, LLMs should interact with users to guide and inspire them to articulate their real intentions.
  • Unclear Table Sources. Existing research often assumes that table sources for answering users’ questions are clearly provided in a single database, spreadsheet, or document. However, in practice, it is frequently unclear which table files are suitable for answering users’ questions. Sometimes, multiple sources of tables are required to address a single query. As a first step, table retrieval might be necessary. Subsequently, merge operations may be required to combine the reasoning results from multiple sources after they have been processed separately. Further reasoning can then only be performed on the merged results.
  • Seperate Models. The mainstream approach involves building separate models for different types of table tasks, particularly for table-related,
    spreadsheet-related, and database-related tasks. While some studies attempt to develop all-in-one large table models for multiple downstream tasks and propose universal prompting frameworks, the distinct characteristics of different table formats and the varying core issues of table tasks in different user scenarios make it highly challenging to find a one-size-fits-all technical solution.

### v3

@@ 5.1 Common Workflow of LLM-powered Agents
Research such as Chain-of-Thought (CoT) [20] and ReAct [81] prompt LLMs iteratively, organizing the reasoning process into multiple intermediate
steps. Thus, LLMs address simpler subproblems step by step and progressively build a coherent response. Agent systems are designed to follow this paradigm and typically include modules such as memory, planning, and action [21]. The memory module stores state or observations from the environment or records past actions. This information can be utilized for future planning. The planning module chooses which action the agent needs to do in the current step, while the action module interacts with the environment and executes the action to get the outcomes. As shown in Figure 8, for table tasks, the agent first observes the table data and user intent. It generates prompts, decomposes complex tasks, plans actions, executes them on the table environment, and then updates the state or observation. This iterative process is repeated until expectations are met. Studies like SheetAgent [38], Chain-of-Table [31], ReAcTable [15], TAPERA [82], and $E^{5}$ [83] follow this workflow.
@@ 5.2.2 Complex Task Decomposition
Inspired by the CoT and least-to-most [84] prompting methods, researchers instruct LLMs to decompose complex table tasks into simpler sub-tasks. DIN-SQL [85] proposes breaking down the NL2SQL task into subtasks. The process involves identifying the relevant tables and columns associated with the query and producing intermediate sub-queries. DEA-SQL [86] and TabSQLify [87] also follow the principle of decomposing tables into smaller, simplifying subtasks and reducing irrelevant information. For web tables, Dater [33] exploits LLMs as decomposers by breaking down huge evidence (a huge table) into sub-evidence (a small table) and decomposing a complex question into simpler sub-questions and getting SQLs.
@@ 5.3.2 Spreadsheet Manipulation and Data Analysis
Spreadsheet manipulation and data analysis are
highly flexible. First, users’ expectations and requirements vary. Second, there is a wide range of operations and software APIs. Third, a complex task may contain multiple operations, leading to dynamic alterations in the table content.
@@ 5.3.3 Multiple Table Tasks
StructGPT [92] is capable of solving table QA, NL2SQL, and knowledge graph QA by developing three types of actions that target web tables, databases, and knowledge graphs. These actions function as a reader, extracting knowledge from the data, which is then used to assist the LLM in its future planning. The actions for web tables include extracting data or column names, whereas for databases, the actions involve extracting table data or metadata. UniDM [36] manages a variety of data manipulation tasks, such as data cleaning, on data lakes through a three-action process. The first action is extracting context information, including table metadata and related records associated with the task, utilizing this information as demonstrations or background knowledge. The second action involves converting this context information into NL texts suitable for LLMs to reason. The final action employs prompt engineering to formulate the desired prompt. TAP4LLM [93] is a pre-processing toolbox to generate table prompts. Its action includes: 1) selecting appropriate rows and columns from the table (called Table Sampling), 2) integrating them with other table data (relevant external knowledge, metadata) (called Table Augmentation), and 3) serializing this information to fit the context length of the LLM. ChatPipe [94] is a system designed to facilitate effortless interaction between users and ChatGPT for table data analysis tasks. Initially, users upload their dataset along with their query. ChatPipe assists in analyzing the dataset and suggests various data analysis and feature engineering operations. Given the numerous potential data operations, ChatPipe employs a reinforcement learning-based method, utilizing a Deep Q-Network model [95] to learn and recommend the most suitable operations.
@@ 4 Excel online forums,
SchemaPile [108]
Column type annotation,
Column population
NL2SQL
Data analysis
github.com
221k DB schemas
1.7M table definitions
@@
synthetically generated 1M tables (png, HTML)
Traditional benchmarks, such as WikiTQ [32], WikiSQL [99], and Spider [42], are already widely used and studied. We will not elaborate on those here but rather focus on new benchmarks. Table 2 presents recently proposed datasets and benchmarks, along with their data sources, sizes. We summarize the features of these new benchmarks as follows:
@@
  • Human Involved Labeling Some new datasets, which target data analysis, require extensive manual annotation. DS-1000 [45], InfiAgent-
    DABench [17], Tapilot-Crossing [97] are designed to assess data analysis tasks. They gather data from the internet and annotate it either automatically or semi-automatically. DS-1000 collects questions from StackOverflow, assesses their usefulness manually, and curates to form the benchmark. The authors manually adapt the original questions by providing input and output context into test cases and rewriting problems to prevent LLMs from learning and memorizing the data. InfiAgent-DABench invites human experts to evaluate the dataset quality and compare human-made and GPT-4 generated data analysis questions via multiple metrics. Tapilot-Crossing aims to build a benchmark for real-world data analysis. An issue that cannot be overlooked is the quality of these datasets. Wretblad et al. conduct a thorough analysis of the NL2SQL dataset BIRD [101], discovering inaccuracies in some of the gold SQLs and noise within certain NL queries [109]. The creators of the BIRD dataset introduce this noise during the dataset creation process. After correcting these errors, Wretblad et al. find that complex prompting methods (e.g., DIN-SQL) might be less effective than simple zero-shot prompting. This research reveals the potential unreliability of table datasets such as BIRD, considering they are generated through human labeling, a process prone to introducing noise and errors.
  • Real-world Workload Most datasets are derived from the web or synthesized using templates. These publicly available online data are usually simple, as some of them are just tutorials for beginners. In contrast, tables in real-world scenarios are far more complex than these. ScienceBenchmark [104] introduces a real-world benchmark developed in collaboration with SQL experts and researchers specializing in policy-making, astrophysics, and cancer research. Given the scarcity of real-world data in these domain-specific databases, ScienceBenchmark employs a data augmentation strategy that starts with hundreds of human-labeled NL2SQL pairs to create thousands more data points. SpreadsheetBench [39] proposes a benchmark aimed at real-world scenarios, where the authors meticulously analyzed user questions from four Excel forums. They argue that their benchmark could assess the performance of handling complex user instructions.
  • Larger Scale AnaMeta [105] is a large-scale table metadata dataset. GitTables [107] downloads millions of CSV tables from GitHub and aligns them with knowledge bases. SchemaPile [108] is a corpus with 221k database schemas and 1.7 million table definitions. These datasets can help evaluate whether the AI system could understand the table schema and benchmark tasks like column type annotation. They can also be fed into neural models as training data for solving various downstream tasks. ComplexTable [46] contains more than 1 million image tables featuring complex structures and can be utilized for visual tasks related to tables.
@@ 6.2 Open-source Software
7 Analysis and Discussion
Table 3: A comparative analysis of various methodologies was conducted using four benchmarks in table QA and NL2SQL tasks. The experimental setup and performance metrics are referenced from Zhang et al. [67]. The number of parameters and training tokens are derived from the respective papers detailing each method.
Type Method # of Parameters # of Tokens to Train Benchmarks Avg. # of Infers
WikiTQ FeTaQA WikiSQL Spider
SLM Training TaPEX 0.14B - 38.55 - 83.90 15.04 1
TaPas 0.11B - 31.60 - 74.20 23.05 1
LLM Training TableLlama 7B 3.3B 48.82 67.73 43.70 - 1
TableLLM 13B 1.1B 62.40 74.50 90.70 83.40 1
Prompting CodeLlama 13B - 43,44 57.24 38.30 21.88 1
GPT-3.5 - - 58.45 71.18 81.70 67.38 1
GPT-4 - - 74.09 78.35 84.00 69.53 1
Prompting Agent StructGPT(GPT-3.5) - - 52.45 11.80 67.80 84.80 3
Binder(GPT-3.5) - - 61.61 12.77 78.60 52.55 50
DATER(GPT-3.5) - - 53.40 18.26 58.20 26.52 100
In order to demonstrate the accuracy and costs of various methods to the readers, this section presents a comparative analysis and discusses the advantages and disadvantages of different approaches.
Specifically, utilizing data from several papers [67, 39, 77, 44, 17, 114], we summarize both accuracy and cost for four table tasks: table QA, NL2SQL, spreadsheet manipulation, and data analysis.
7.1 Discussion on LLM Training
The advantages of training-based methods are that they offer great control, allowing enterprises to conduct private training and deployment without data leakage to third-party model service providers. However, two issues can’t be ignored: cost and accuracy.
Cost The expense of pre-training or fine-tuning a large model is substantial. Fine-tuning a 7B model typically requires eight 80GB GPUs, and the number of training tokens dictates the total time needed for training. The training speed, measured in tokens per second, is affected by both the hardware and the software and is reported in several technical reports, such as Llama [115]. Table 3 lists the number of parameters and training tokens for two LLM training approaches, thereby providing readers with an understanding of the training costs. The number of parameters can also be used to estimate the minimal costs for private deployment. For instance, loading a 7B model with float16 format requires at least 14GB of memory, while serving with concurrent requests necessitates additional memory for the transformer architecture’s key-value (KV) cache [116].
Accuracy As indicated in Table 3, across four benchmarks, the LLM training approach excels in only one, whereas prompting a robust LLM (GPT-4) is superior. The 7B TableLlama can’t outperform task-specific fine-tuning models on specific tasks (see the WiKiSQL column in Table 3). Another example is the NL2SQL task, where Li et al. [77] perform a systematic evaluation with various methods, comparing LLM-based and SLM-based solutions. Results show that there is no clear winner between LLM and SLM solutions on different metrics and domains. On data analysis tasks, instruction tuned models like Lemur [44] and DAAgent [17] cannot surpass GPT-4. Table-GPT continues to train on GPT-3.5 and achieves better results on all table-related tasks than GPT-3.5 and ChatGPT. However, the cost of training is
prohibitively high for ordinary enterprise users who wish to deploy privately. Regarding training data, the cost of manual annotation is also high. Although synthesis is a cheap choice, the data quality is another concern. A more prevalent approach is using GPT as a teacher model for data distillation. Evidently, GPT is the performance ceiling.
7.2 Discussion on LLM Prompting
The LLM-powered agent method combines the
power of LLMs and the flexibility of external tools. On many table tasks, prompting strong LLMs like GPT still performs the best.
Cost As shown in the “Avg. # of Infers” column of Table 3, agents must prompt LLMs repeatedly to fulfill users’ expectations. Ma et al. [39] conduct experiments on agents for spreadsheet manipulation, and the most effective agent solution is the one that utilizes the ReAct framework [81], executes code within the execution environment and offers feedback to LLMs upon failure. Deploying multiple rounds of ReAct prompting with execution feedback takes a substantial time. Consequently, all these factors increase the total time and financial costs.
Structure Understanding Foundation models, without fine-tuning for tables, still have difficulty understanding table structure and hierarchy. Pang et al.  [114] create a benchmark named TIS to assess how effectively LLMs seek information from tables. They discover that LLMs struggle with tables that have complex hierarchies and exhibit a poor understanding of table structures, such as locating a specific regin in a two-dimensional table. Notably, most LLMs perform at nearly random accuracy levels (around 50%) in the table structure understanding task, while GPT-4 achieve 66.1%.
@@ 8 Challenges and Future Directions
8.1 Diverse User Input When Serving
User inputs refer not only to users’ NL queries but also to the table schema and content.
User Query In real-world applications, user NL queries are frequently ambiguous. For example, when users are unfamiliar with the table they want to query, they often pose general requests without a clear objective, such as “help me analyze this table”. Users might also pose questions unrelated to the table within a table AI system. Even when users have some knowledge of the table schema and content, they may struggle to formulate their query accurately, leading to ambiguity in the query sentences.
Table Schema and Content In practice, table schemas and content are highly diverse and often proprietary. For instance, in a domain or industry with which the model has limited knowledge, it faces challenges in effectively transferring into the field. Existing table training datasets are relatively simple, either consisting of simple table structures scraped from the web or synthesized based on powerful LLMs. Constructing a complex and diverse training dataset is quite costly. Ensuring that the training data covers real-world business scenarios poses a significant challenge.
Future table LLMs should adapt quickly and cheaply to real-world business needs. Research directions include synthesizing high-quality training data that reflects the diverse needs of specific domains by cost-effective methods.
8.2 Slow and Deep Thinking
Daniel Kahneman has revealed that slow and deep thinking is the underlying mechanism of the human brain for processing complex problems [117]. The chain-of-thought approach is considered to guide LLMs in solving complex reasoning problems, and the recent OpenAI’s o1 shows that LLMs can achieve substantial enhancements in performance by allowing them to generate extended internal chains-of-thought. Additionally, this internal chain-of-thought can facilitate both the model training processes and the pure model prompting methods.
Training Typically, there is a considerable gap between the user instructions and the ultimate answer. Therefore, a chain-of-thought can serve as an internal mechanism to explain the process of deriving the final answer, potentially mitigating the training difficulty faced by LLMs. However, acquiring an accurate chain-of-thought is challenging. It can be prohibitively expensive when achieved through human annotation and difficult to ensure accuracy when automatically generated by LLMs themselves, particularly in tasks involving complex logical reasoning over tables. Exploring cost-effective methods to guarantee an accurate chain-of-thought process is a valuable area of research.
Prompting In existing prompting methods on table tasks, researchers develop their own chain-of-thought workflows, which heavily rely on hard-coded prompt templates to solve specific table tasks. This results in methods with weak transferability. Given that OpenAI’s o1 possesses inherent chain-of-thought abilities, prompting and agent methods must reconsider how to construct their workflows. On the other hand, the chain-of-thought approach requires numerous inference iterations, which is time-consuming; future table prompting methods should balance inference times with accuracy.